1.5x+3=23
2.4=3a-14
3.19=3y-5

Answer by deborabr(181)

you just have group the single numbers with others single numbers and numbers unknown with numbers unknown but they need be equal, for example, "x" only with "x", "a" only with "a" and "y" only with "y". but dont forget of to change the signal when you pass of one side to other and if this number is multiplying pass it dividing and vice-versa, like this is how we isolate the unknown and find the values
1.5x+3=23
5x=23-3
5x=20
x=20/5
x=4
2.4=3a-14
4+14=3a
18=3a
18/3=a
6=a
3.19=3y-5
19+5=3y
24=3y
24/3=y
8=y
Answers:
1. x=4
2. a=6
3. y=8